Given the formula for change in internal energy, U = Q + W, calculate work for a



system that absorbs 350 kJ of heat and for which U = 150 kJ. Is the work done on



the system or by the system? Does the system expand or contract?

1
Expert's answer
2022-04-20T03:57:03-0400

W = U - Q = 150 - 350 = -200 kJ

Work is negative, which means work is done by the system on its surroundings.

W = -pdV, W is negative, then dV is positive, which means that gas expands.